# Generalization in Reinforcement Learning for Radio Access Networks

### v2

@@ 7.4 State Space Design
To enable distributed learning of a single RL policy from the collective experience of all UEs across the network, we model the state of the LA process by combining (a) a UE-centric view of the RAN environment surrounding the UE, consisting of semi-static information characterizing network nodes, radio cells, and the UE itself (see Section 4), and (b) dynamic information characterizing relevant features of the LA process. Together with domain randomization during training (see Section 3), this allows us to train models that generalize across the RAN environment.
@@ 7.4.1 Semi-static Information
We describe the network deployment surrounding a UE by considering its serving cell and a subset of the most relevant interfering cells, as illustrated in Figure 1. Specifically, we select $K_{I}$ cells whose signal strength, measured by the UE, exceeds a predefined threshold. Thus, the composition of surrounding cells may change over time depending on UE mobility. For each cell, we include semi-static configuration information such as site type (e.g., LTE or NR), carrier frequency, antenna array configuration (e.g., MIMO or mMIMO), geometric properties (e.g., inter-site distance), bandwidth, and transmit power.
Since the strength and quality of signals perceived by a UE can vary significantly due to manufacturing characteristics—such as chipset, antenna hardware, and algorithms for signal processing or channel state estimation—the MDP state should capture information specific to each UE to generalize effectively across the diversity of the UE population. To this end, we include indicators representing the UE type and antenna array size.
In our design, we provide this information either directly to a DQN or graph-based model, as described in Section 4. While our numerical evaluations (see Section 8) rely on a limited set of semi-static parameters sufficient for characterizing typical RAN deployments, additional parameters may be beneficial for practical deployment in live networks.
@@ 7.4.2 Dynamic Information
We consider dynamic, observable information relevant to the LA process for a UE before packet transmission, encompassing measurements of radio link conditions, resource availability, and the UE’s state. As LA operates on a sub-millisecond timescale, accounting for information aging is crucial to accurately predict the optimal MCS index. We thus apply either a soft or hard aging model to different information types, as detailed below.
Packet Transmission Counter: Since retransmissions can take advantage of previously received copies of the packet to improve the decoding probabilities, we include the packet transmission counter $n$ (see Equation 2) among the state features. Differentiating between initial transmissions and retransmissions allows policies to effectively optimize LA parameters in multiple transmission attempts of the same packet.
@@
Historical CSI: Instead of relying solely on the most recent CSI report, we consider a sequence of $K_{CSI}\geq 1$ historical CQIs and rank values, along with a rank-weighted average CQI. To account for information aging, we define a time window $W$, setting CQI and rank values to $-1$ for reports older than $W$. The window $W$ can correspond to the channel coherence time, allowing models to prioritize recent channel information when determining LA parameters.
UE Buffer Status: The UE’s buffer status indicates the amount of data pending transmission. Instead of using the raw buffer size $B_{t}$ at time $t\in\mathbb{R}$, we normalize $B_{t}$ by the maximum transport block size $TBS_{\max}$ allowed by the system:
@@
where $SE_{k}^{(i)}$ and $rank_{k}^{(i)}$ denote the spectral efficiency and transmission rank associated with the $k$-th historical CSI report for UE $i$. Specifically, $SE_{k}^{(i)}$ corresponds to the nominal spectral efficiency associated with the $k$-th reported CQI, as defined in the 3GPP 5G NR specifications 3rd Generation Partnership Project (3GPP).
@@
Historical LA Parameters: Historical LA parameters (MCS index, rank) from previous transmissions also aid state reconstruction. We include $K_{LA}$ historical MCS indices $\{m_{k}^{(t_{k})}\}_{k=1}^{K_{LA}}$ and rank values $\{l_{k}^{(t_{k})}\}_{k=1}^{K_{LA}}$. To account for aging, we employ a hard aging model, assigning a value of $-1$ for historical data older than maximum age $W$:
@@ 8.1.1 Distributed Training Architecture
Our training system, illustrated in Figure 2, utilizes a distributed training architecture comprising a single learner, 40 actors, and a shared replay memory composed of several independent shards. These components are co-located on a shared compute node, while the simulation environments are hosted on separate compute nodes to facilitate distributed and scalable experience collection. To maximize throughput, the system operates asynchronously, allowing actors and the learner to work independently without global synchronization.
The replay memory is partitioned into four independently prioritized shards, each maintaining its own priority queue. Each shard has a capacity of 4 million transitions, enabling long-term experience storage and sampling diversity. Actors write transitions asynchronously to specific shards in batches, reducing communication overhead and supporting scalable data ingestion, similar to Kapturowski et al. (2022).
As shown in Figure 5, each actor steadily writes batches to its assigned shard at a rate of approximately 18–20 batches per minute after an initial ramp-up period. This time series, covering a representative 30-hour training run, displays data from four actors (IDs 0, 10, 20, and 30) and highlights both the system’s ability to reach a stable ingestion regime and the even distribution of load across actors and shards. The close alignment in write rates across disparate actor indices provides strong empirical evidence of the scalability and load-balancing capabilities of our distributed experience collection infrastructure.
@@
Actors run local replicas of the policy network and interact concurrently with 14 environment instances (see Section 10.2 that provides additional computation and deployment details). Similar to Horgan et al. (2018), each actor $i\in\{0,\cdots,39\}$ executes an $\epsilon_{i}$-greedy policy, defined as $\epsilon_{i}=\epsilon^{1+\frac{i}{N-1}\alpha}$, where $\epsilon$ is set to 0.4 and $\alpha$ to 8.5. This exploration policy remains fixed throughout training. On average, each actor generates approximately 132 transitions per second. Incoming transitions are buffered locally before being dispatched asynchronously to replay memory in batches of 500 transitions. Prior to training, a warm-up phase collects 45,000 transitions—without gradient updates—to populate the replay memory with a diverse initial dataset.
The learner is responsible for computing gradients and updating model parameters. It prefetches up to 16 batches of 512 transitions each from the replay memory shards. This prefetching ensures a continuous supply of data for gradient computation. On average, the learner processes 100 batches per second, corresponding to approximately 51,200 transitions per second. Model parameters are updated and broadcast asynchronously to all actors every 512 learner steps (roughly every 5 seconds), allowing actors to operate with relatively fresh policies while maintaining a decoupled learning and acting process.
@@
Carrier frequency 3.5 GHz Fixed
@@
Cell radius {166, 300, 600, 900, 1200} m Randomized
Bandwidth {20, 40, 50, 80, 100} MHz Randomized
@@
DL TX power {20, 40, 50, 80, 100} W Randomized
@@
Speed UE FB {0.67, 10, 15, 30} m/s Randomized
Speed UE eMBB {0.67, 1.5, 3} m/s Randomized
@@ 8.1.2 Training Environment Diversification
We train the RL policy to generalize in the RAN environment by exposing it to a wide range of deployment scenarios, radio conditions, traffic patterns, interference levels, and load distributions, as discussed in Section 3.2. While such diversity is naturally present in live RAN systems—where training data can be generated across network nodes—we use domain randomization to recreate comparable conditions within network simulators.
@@
Each training scenario is generated using an event-driven, 5G-compliant system simulator that operates in time division duplexing (TDD) mode on a 3.5 GHz carrier frequency, with single-user multiple input multiple output (SU-MIMO) transmission and a physical layer model based on numerology $\mu=0$ as specified in 3GPP NR standard 38.211 3rd Generation Partnership Project (3GPP).
A training scenario consists of three trisectorial radio sites, each randomly configured as either MIMO or mMIMO. Site configurations are randomized with respect to cell radius, system bandwidth, and downlink transmit power, using parameters sampled from Table 1. Further randomization is applied to cell load, traffic type, indoor/outdoor UE distribution, andUE receiver types. A UE generation process randomly selects the number of UEs with full buffer (FB) and eMBB traffic to be deployed, drawing from a predefined set of indoor/outdoor probabilities listed in Table 1.
Each eMBB UE generates traffic based on a packet arrival process and a packet size distribution derived from historical field data, with payloads ranging from 20 to approximately 1000 Kbytes. Each UE configuration is then randomized by the number of antennas, mobility speed, and receiver type. The latter emulates variations in channel estimation accuracy associated with differentUE chipsets under the same channel conditions.
For hyperparameter tuning, we train RL policies on 40,000 training scenarios generated from 160 random seeds and 250 randomized simulation parameter configurations, as defined in Table 1. Each scenario simulates five seconds (equivalent to 5,000 transmission time intervals), producing approximately 16,000 training samples on average and producing a total of approximately 650 million samples across the entire sweep. The resulting hyperparameters are summerized in Section 10.1.
For the general performance evaluations presented in Section 8.3, we employ a larger dataset of 240,000 training scenarios generated from the same 160 seeds but with 1,500 randomized parameter configurations, also drawn from Table 1. This setup produces approximately 3 billion training samples in total.
@@
Figure 6: Performance comparison of specialized and generalized RL-based LA policies across five benchmark scenarios (B1–B5). The boxplots illustrate relative gain or loss (%) over a state-of-the-art baseline for two RL-based LA policies, both employing the same neural network architectures—a six-layer multilayer perceptron with 256 neurons per layer. The policies differ only by training procedure: a specialized policy trained exclusively for full-buffer (FB) conditions (red), and a generalized policy trained across varied network conditions (blue). Results are presented for five benchmark scenarios (B1–B5), showing the distribution of relative performance gains in: (a) downlink cell throughput, and (b) spectral efficiency. Each boxplot displays the median (diamond), interquartile ranges (box), whiskers extending to 2×2$\times$2 ×IQR, and outliers (circles).
@@ 8.2 Test Setup
For testing purposes, we define five benchmark scenarios, described in Table 2, whose conditions are not directly encountered during training. Benchmark B1 consists of a single-cell, single-UE deployment positioned randomly and operating in MIMO transmission mode. This scenario represents a challenging outlier for evaluating the link-curve performance of RL policies against the state-of-the-art OLLA algorithm used in 5G networks, which is specifically optimized for such conditions.
Scenarios B2–B5 consist of homogeneous network deployments—either MIMO or mMIMO—with simulation parameters based on the default values indicated in bold in Table 1. Each benchmark targets specific testing conditions: B2 evaluates stable interference under full-buffer traffic using either MIMO (B2(a)) or mMIMO (B2(b)) deployments; B3 simulates dynamic interference in mMIMO deployments due to eMBB traffic; B4 assesses performance for UEs traveling at high speeds (100–200 km/h); and B5 evaluates performance under realistic traffic conditions involving a mix of FB file transfers and chat traffic types not used during training.
@@ 8.3 General Results
To validate our design concepts for RL generalization in RAN applications, we compare the performance of two LA policies trained using the same model architecture and hyperparameters but under different data regimes. The first policy is trained to generalize using randomized network scenarios, while the second is specialized for FB traffic conditions by using an equal amount of simulated networks based on benchmarks B2(a), B2(b), and B4, with randomized seeds. In both cases, we use an MLP with six hidden layers and 256 neurons per layer. Inputs are described in Section 7.4, and the output layer comprises 28 nodes, each representing the Q-value associated with an MCS index compliant with 5G specifications (Table 5.1.3.1-2 in 3GPP TS 38.214 3rd Generation Partnership Project (3GPP)). Both policies are trained using a robustness weight $\alpha=0.5$ (see Equation 2).
We compare both policies with a state-of-the-art outer-loop LA design that targets a $10\%$ BLER, representing the most common approach adopted in real-world 5G NR systems.
Figure 6 summarizes the test results obtained across the five benchmarks described in Table 2, reporting statistics of gains and losses in average cell throughput (see Figure 6) and spectral efficiency (see Figure 6) relative to the baseline.
First, we observe that the generalized RL policy significantly outperforms the 5G baseline across all benchmarks, achieving gains in average cell throughput and spectral efficiency close to or exceeding 10% in FB MIMO and mMIMO scenarios, and surpassing 20% in high-mobility conditions.
@@
Additional evaluations on hyperparameter tuning—including model architecture, layer normalization, and loss functions—are presented in Section 10.3.
@@
Figure 7: Performance analysis in single-cell single-UE benchmark B1, with the model tested on several realization of the scenario with different UE positions ranging from 40 to 1250 meters from the serving cell. Throughput gains are measured with respect to the OLLA baseline.
@@ 8.3.1 Link-Level Performance
Figure 7 further examines the performance of both models in the single-cell single-UE benchmark B1. This benchmark is relevant for two main reasons: it reflects outlier conditions never encountered during training, and it represents the scenario for which the OLLA baseline is optimized. The latter provides reference values for link-level spectral efficiency used, for instance, by 3GPP to standardize CQI and MCS tables 3rd Generation Partnership Project (3GPP).
We evaluate both the generalized and specialized models against the baseline across several initial UE locations, ranging from 40 meters to 1250 meters from the serving base station. At each location, we conduct 50 simulations using distinct random seeds for channel and environment realizations. Figure 7 shows that the generalized RL model consistently outperforms both the baseline and the specialized model at all locations. While average throughput gains range between 2% and 5.5%, it is important to note that the OLLA baseline nearly reaches link capacity under B1 conditions, leaving limited room for improvement.
Figure 7 shows that the generalized RL model achieves these gains by selecting, on average, slightly more conservative MCS indices at greater distances from the base station. This conservatism reduces the number of retransmissions experienced by the RL policy, resulting in modest but consistent throughput gains. In contrast, the specialized model is overly conservative: near the base station, it fails to recognize ideal channel conditions (i.e., low path loss attenuation), instead selecting lower MCS indices. This results in fewer information bits being transmitted and thus a performance loss. At other distances, this model also tends to select slightly more conservative MCS indices, yielding smaller gains over the baseline.

### v3

@@ 8.4 State Space Design
To enable distributed learning of a single RL policy from the collective experience of all UEs across the network, we model the state of the LA process by combining (a) a UE-centric view of the RAN environment surrounding the UE, consisting of semi-static information characterizing network nodes, radio cells, and the UE itself (see Section˜4), and (b) dynamic information characterizing relevant features of the LA process. Together with domain randomization during training (see Section˜3), this allows us to train models that generalize across the RAN environment.
@@ 8.4.1 Semi-static Information
We describe the network deployment surrounding a UE by considering its serving cell and a subset of the most relevant interfering cells, as illustrated in Figure˜1. Specifically, we select $K_{I}$ cells whose signal strength, measured by the UE, exceeds a predefined threshold. Thus, the composition of surrounding cells may change over time depending on UE mobility. For each cell, we include semi-static configuration information such as site type (e.g., LTE or NR), carrier frequency, antenna array configuration (e.g., MIMO or mMIMO), geometric properties (e.g., inter-site distance), bandwidth, and transmit power.
Since the strength and quality of signals perceived by a UE can vary significantly due to manufacturing characteristics—such as chipset, antenna hardware, and algorithms for signal processing or channel state estimation—the MDP state should capture information characterizing different UE types to generalize effectively across the diversity of the UE population. To this end, we include indicators representing the UE type and antenna array size.
We feed this information to either a DQN or a graph-based model, as described in Section˜4. While a limited set of semi-static parameters is sufficient to model the RAN deployment in our simulations (cf. Section˜9), additional parameters may be beneficial to better characterize the diversity of real-world network deployments.
@@ 8.4.2 Dynamic Information
We consider dynamic, observable information relevant to the LA process for a UE before a packet transmission, encompassing measurements of radio link conditions, resource availability, and the UE’s state. As LA operates on a sub-millisecond timescale, accounting for information aging is crucial to accurately predict the optimal MCS index. We thus apply either a soft or hard aging model to different information types, as detailed below.
Packet Transmission Counter: Since retransmissions can take advantage of previously received copies of the packet to improve the decoding probability, we include a packet transmission counter $n$ (see Equation˜2) among the state features. Differentiating between the initial transmission and subsequent retransmissions allows to learn policies that effectively optimize LA parameters in multiple transmission attempts of the same packet.
@@
Historical CSI: Instead of relying solely on the most recent CSI report, we consider a sequence of $K_{CSI}\geq 1$ historical CQIs and rank values, along with a rank-weighted average CQI. To account for information aging, we define a time window $W$, setting CQI and rank values to 1-1 for reports older than $W$. The window $W$ can correspond to the channel coherence time, allowing models to prioritize recent channel information when determining LA parameters.
UE Buffer Status: The UE buffer status indicates the amount of data pending transmission. Instead of using the raw buffer size $B_{t}\in\mathbb{R}_{+}$ at time $t\in\mathbb{N}$, we normalize $B_{t}$ by the maximum transport block size $TBS_{\max}$ allowed by the system:
@@
where $SE_{k}^{(i)}$ and $rank_{k}^{(i)}$ denote the spectral efficiency and transmission rank associated with the $k$-th historical CSI report for UE $i$. Specifically, $SE_{k}^{(i)}$ corresponds to the nominal spectral efficiency associated with the $k$-th reported CQI, as defined in the 3GPP 5G NR specifications 3GPP38214.
@@
Historical LA Parameters: Historical LA parameters (MCS index, rank) from previous transmissions also aid state reconstruction. We include $K_{LA}$ historical MCS indices $\{m_{k}^{(t_{k})}\}_{k=1}^{K_{LA}}$ and rank values $\{l_{k}^{(t_{k})}\}_{k=1}^{K_{LA}}$. To account for aging, we employ a hard aging model, assigning a value of 1-1 for historical data older than maximum age $W$:
@@ 9.1.1 Distributed Training Architecture
Our training system, illustrated in Figure˜2, utilizes a distributed architecture comprising a single learner, 40 actors, and a shared replay memory composed of several independent shards. These components are co-located on a shared compute node, while the simulation environments are hosted on separate compute nodes to facilitate a distributed and scalable experience collection. To maximize throughput, the system operates asynchronously, allowing actors and the learner to work independently without global synchronization.
The replay memory is partitioned into four independently prioritized shards, each maintaining its own priority queue. Each shard has a capacity of 4 million transitions, enabling long-term experience storage and sampling diversity. Actors write transitions asynchronously to specific shards in batches, reducing communication overhead and supporting scalable data ingestion, similar to KCJ+:22.
As shown in Figure˜5, each actor steadily writes batches to its assigned shard at a rate of approximately 18–20 batches per minute after an initial ramp-up period. This time series, covering a representative 30-hour training run, displays data from four actors (IDs 0, 10, 20, and 30) and highlights both the system’s ability to reach a stable ingestion regime and the even distribution of load across actors and shards. The close alignment in write rates across disparate actor indices provides strong empirical evidence of the scalability and load-balancing capabilities of our distributed experience collection infrastructure.
@@
Actors run local replicas of the policy network and interact concurrently with 14 environment instances (Section˜11.2 provides additional computation and deployment details). Similar to Horgan:2018, each actor $i\in\{0,\cdots,39\}$ executes an $\epsilon_{i}$-greedy policy, defined as $\epsilon_{i}=\epsilon^{1+\frac{i}{N-1}\alpha}$, where $\epsilon$ is set to 0.4 and $\alpha$ to 8.5. This exploration policy remains fixed throughout training. On average, each actor generates approximately 132 transitions per second. Incoming transitions are buffered locally before being dispatched asynchronously to a replay memory in batches of 500 transitions. Prior to training, a warm-up phase collects 45,000 transitions—without gradient updates—to populate the replay memory with a diverse initial dataset.
[…]
@@
Carrier frequency 3.5 $\mathrm{GHz}\text{/}$ Fixed
@@
Cell radius {166, 300, 600, 900, 1200} $\mathrm{m}\text{/}$ Randomized
Bandwidth {20, 40, 50, 80, 100} $\mathrm{MHz}\text{/}$ Randomized
@@
DL TX power {20, 40, 50, 80, 100} $\mathrm{W}\text{/}$ Randomized
@@
Speed UE FB {0.67, 10, 15, 30} $\mathrm{m}\text{/}\mathrm{s}$ Randomized
Speed UE eMBB {0.67, 1.5, 3} $\mathrm{m}\text{/}\mathrm{s}$ Randomized
@@ 9.1.2 Training Environment Diversification
We train the RL policy to generalize in the RAN environment by exposing it to a wide range of deployment scenarios, radio conditions, traffic patterns, interference levels, and load distributions, as discussed in Section˜3.2. While such diversity is naturally present in live RAN systems—where training data can be generated across network nodes—we use domain randomization to recreate comparable conditions within network simulators.
@@
[…]
A training scenario consists of three trisectorial radio sites, each randomly configured as either MIMO or mMIMO. Site configurations are randomized with respect to cell radius, system bandwidth, and downlink transmit power, using parameters sampled from Table˜1. Further randomization is applied to cell load, traffic type, indoor/outdoor UE distribution, and UE receiver types. A UE generation process randomly selects the number of UEs with full buffer (FB) and eMBB traffic to be deployed, drawing from a predefined set of indoor/outdoor probabilities listed in Table˜1.
[…]
Each UE configuration is further randomized by the number of antennas, mobility speed, and receiver type. The latter accounts for variations in channel estimation accuracy among different UE chipsets under identical channel conditions.
For hyperparameter tuning, summarized in Section˜11.1, we train RL policies on 40,000 training scenarios generated from 160 random seeds and 250 randomized simulation parameter configurations, as defined in Table˜1. Each scenario simulates five seconds (equivalent to 5,000 transmission time intervals), producing approximately 16,000 training samples on average and producing a total of approximately 650 million samples across the entire sweep.
Section˜9.3 presents the general performance evaluation employing a larger dataset of 240,000 training scenarios generated from the same 160 seeds but with 1,500 randomized parameter configurations, also drawn from Table˜1. This setup produces approximately 3 billion training samples.
@@
Figure 6: Performance comparison of specialized and generalized RL-based LA policies across five benchmark scenarios (B1–B5). The boxplots illustrate relative gain or loss (%) over a state-of-the-art baseline for two RL-based LA policies, both employing the same neural network architectures: a six-layer multilayer perceptron with 256 neurons per layer. The policies differ only by training procedure: a specialized policy trained exclusively for full-buffer (FB) conditions (red), and a generalized policy trained across varied network conditions (blue). Results are presented for five benchmark scenarios (B1–B5), showing the distribution of relative performance gains in: (a) downlink cell throughput, and (b) spectral efficiency. Each boxplot displays the median (diamond), interquartile ranges (box), whiskers extending to $2\timesIQR$, and outliers (circles).
@@ 9.2 Test Setup
For testing purposes, we define five benchmark scenarios, summarized in Table˜2, whose conditions are not directly encountered during training. Benchmark B1 corresponds to a single-cell, single-UE deployment operating in MIMO transmission mode. This benchmark is used to assess link-level performance, where the UE is placed at predefined distances from the base station, and the channel realization is randomized across multiple simulations per UE position. This scenario serves as a challenging outlier for evaluating the link-level performance of RL-based policies against the state-of-the-art OLLA algorithm used in 5G networks, which is specifically optimized for such conditions.
[…]
@@ 9.3 General Results
To validate our design concepts for RL generalization in RAN applications, we compare the performance of two LA policies trained using the same model architecture and hyperparameters but under different data regimes. The first policy is trained to generalize using randomized network scenarios, while the second is specialized for FB traffic conditions by using an equal amount of simulated networks based on benchmarks B2(a), B2(b), and B4, with randomized seeds. In both cases, we use an MLP with six hidden layers and 256 neurons per layer. Inputs are described in Section˜8.4, and the output layer comprises 28 nodes, each representing the Q-value associated with an MCS index compliant with 5G specifications (Table 5.1.3.1-2 in 3GPP TS 38.214 3GPP38214). Both policies are trained using a robustness weight $\alpha=0.5$ (see Equation˜2).
We compare both policies with a state-of-the-art OLLA design that targets a $10\%$ BLER, representing the most common approach adopted in real-world 5G NR systems.
Figure˜6 summarizes the test results obtained across the five benchmarks described in Table˜2, reporting statistics of gains and losses in average cell throughput (cf. Figure˜6) and spectral efficiency (cf. Figure˜6) relative to the baseline.
First, we observe that the generalized RL policy significantly outperforms the 5G baseline across all benchmarks, achieving gains in average cell throughput and spectral efficiency close to or exceeding 10% in FB MIMO and mMIMO scenarios (i.e., B2(a)-(b)), and approximately 20% in high-mobility conditions (i.e., B4).
@@
Additional evaluations on hyperparameter tuning, including model architecture, layer normalization, and loss functions, are presented in Section˜11.3.
@@
Figure 7: Performance analysis in single-cell single-UE benchmark B1, with the model tested on several realizations of the scenario with different UE positions ranging from 40 m to 1250 m from the serving cell. Throughput gains are measured with respect to the OLLA baseline.
@@ 9.3.1 Link-Level Performance
Figure˜7 further examines the performance of both models in the single-cell single-UE benchmark B1. This benchmark is relevant for two main reasons: it reflects outlier conditions never encountered during training, and it represents the scenario for which the OLLA baseline is optimized. The latter provides reference values for link-level spectral efficiency used, for instance, by 3GPP to standardize CQI and MCS tables 3GPP38214.
[…]
Figure˜7 shows that the generalized RL model achieves these gains by selecting, on average, slightly more conservative MCS indices at greater distances from the base station. This conservatism reduces the number of retransmissions experienced by the RL policy, resulting in modest but consistent throughput gains. In contrast, the specialized model is overly conservative: near the base station, it fails to recognize ideal channel conditions (i.e., low path loss attenuation), selecting instead lower MCS indices. This results in fewer information bits being transmitted and thus a performance loss. At other distances, this model also tends to select slightly more conservative MCS indices, yielding smaller gains over the baseline.